OpenStudy (anonymous):

integral of ((3x-2)^2 )/(x)^1/2

OpenStudy (anonymous):

just multiply out the numerator then simplify. the integral should be easier then....

OpenStudy (anonymous):

\[\frac{(3x - 2)^{2}}{x^{\frac{1}{2}}}=\frac{9x^{2}-12x+4}{x^{\frac{1}{2}}}\]\[=9x^{\frac{3}{2}}-12x^{\frac{1}{2}}+4x^{-\frac{1}{2}}\]